### v3

# VLCE: A Knowledge-Enhanced Framework for Image Description in Disaster Assessment

Abstract
Immediate damage assessment is essential after natural catastrophes; yet, conventional hand evaluation techniques are sluggish and perilous. Although satellite and unmanned aerial vehicle (UAV) photos offer extensive perspectives of impacted regions, current computer vision methodologies generally yield just classification labels or segmentation masks, so constraining their capacity to deliver a thorough situational comprehension. We introduce the Vision Language Caption Enhancer (VLCE), a multimodal system designed to produce comprehensive, contextually-informed explanations of disaster imagery. VLCE employs a dual-architecture approach: a CNN-LSTM model with a ResNet50 backbone pretrained on EuroSat satellite imagery for the xBD dataset, and a Vision Transformer (ViT) model pretrained on UAV pictures for the RescueNet dataset. Both systems utilize external semantic knowledge from ConceptNet and WordNet to expand vocabulary coverage and improve description accuracy. We assess VLCE in comparison to leading vision-language models (LLaVA and QwenVL) utilizing CLIPScore for semantic alignment and InfoMetIC for caption informativeness. Experimental findings indicate that VLCE markedly surpasses baseline models, attaining a maximum of 95.33% on InfoMetIC while preserving competitive semantic alignment. Our dual-architecture system demonstrates significant potential for improving disaster damage assessment by automating the production of actionable, information-dense descriptions from satellite and drone photos.
Keywords: Disaster Image Captioning; Vision-Language Models; Knowledge Graph Augmentation; Multimodal Deep Learning; Remote Sensing Analysis; Emergency Response Systems; Semantic Knowledge Integration
1 Introduction
Global populations are particularly vulnerable to natural catastrophes, since floods, earthquakes, hurricanes, and wildfires cause extensive harm to ecosystems, infrastructure, and human lives all across the world. In these circumstances, quick damage assessment is essential for directing long-term restoration activities, assisting with search and rescue missions, and assigning emergency resources. However, in the immediate aftermath of disasters, typical manual assessment procedures are sometimes impracticable, slow, and dangerous.
Recent advancements in computer vision and remote sensing technology present novel prospects for disaster damage evaluation. Satellite and unmanned aerial vehicle (UAV) imagery offer high-resolution perspectives of impacted regions, whereas deep learning frameworks have demonstrated significant efficacy in identifying and segmenting damage from aerial photographs [11]. Nonetheless, these techniques generally yield merely class labels or segmentation masks, constraining their capacity to deliver the detailed descriptive information essential for comprehending intricate disaster scenarios.
To alleviate these challenges, image captioning has arisen as an effective method for visual interpretation and contextual comprehension. Image captioning is essential for assessing visual input and producing contextual descriptions that offer semantic information beyond rapid observation. Although photographs proficiently convey visual details and spatial relationships, they may not clearly express temporal context, causal links, or domain-specific interpretations. Image captioning augments visual information by transforming observations into descriptive, comprehensible text, so improving accessibility, interpretability, and actionability for decision-makers.
Example (Disaster Scenario):
Figure 1: RescueNet disaster image of post-hurricane devastation contrasting basic visual description with detailed automatic caption generation.
Without caption: “A satellite image depicts a community from an aerial perspective, revealing multiple dwellings and streets”
With caption: “The image shows the aftermath of Hurricane Michael, which inflicted major damage to infrastructure and the environment. Roads and streets are cluttered with debris from destroyed structures, such as fallen trees and scattered items. Trees in the region have suffered varied degrees of damage, with some seeming damaged or totally fallen. Buildings show different degrees of destruction, ranging from comparatively intact ones to those with obvious traces of catastrophic damage. The entire scene depicts the ongoing recovery and rebuilding activities in the aftermath of the catastrophe.”
This contrast demonstrates how automated captioning goes beyond simple visual description to offer contextual knowledge. Responders may now make better decisions because the improved caption makes it apparent what kind of disaster it is, what infrastructure is impacted, and how far along the recovery is. As a result, image captioning goes beyond simple visual description to offer situational awareness and useful intelligence that is essential for applications involving accessibility, surveillance, and disaster response.
Prior investigations into disaster image analysis have predominantly concentrated on remote sensing methodologies for damage assessment and disaster categorization. Many research have employed satellite image segmentation and change detection techniques [10], although these methods fail to produce descriptive text summaries necessary for thorough situational analysis. Object detection in disaster imagery presents challenges across diverse settings [7], and the precise identification of disaster-affected regions from aerial photography offers substantial difficulty for machine learning models [9]. Multimodal approaches signify sophisticated methodologies for disaster assessment [12], with models such as LLaVA[13] showcasing notable progress in vision-language integration.
Vision-Language Models (VLMs) have recently revolutionized automated image captioning capabilities [5]. These models show significant generalization across various domains; however, they lack proficiency in specialized tasks, underscoring the disparity between generalist capabilities and domain-specific expertise [6]. In critical scenarios like disasters, precise, detailed, and contextually relevant captions are vital for effective situational comprehension [14]. Image captioning in disaster contexts poses distinct challenges stemming from intricate environments, ambiguous object boundaries, and specialized content requirements. Although general-purpose VLMs may exhibit enhanced overall performance, they frequently lack the specialized knowledge required for effective interpretation of disaster scenes.
To overcome these limitations, we propose VLCE (Vision Language Caption Enhancer), a specialized image captioning framework that enhances caption generation by integrating knowledge for the analysis of disaster scenarios. This approach combines visual features with external semantic knowledge from ConceptNet and WordNet to improve semantic understanding and produce descriptive, contextually relevant captions. Our model utilizes cross-modal attention mechanisms within tailored architectures to connect raw visual signals with significant, domain-specific textual descriptions. This framework specifically targets the generation of descriptions for satellite and drone imagery in disaster-affected regions, thereby addressing a notable deficiency in automated disaster assessment capabilities.
The key contributions of this research include:
  • Creation of an innovative dual-architecture framework employing cross-modal attention processes for catastrophe picture captioning, featuring dedicated models for satellite and UAV imagery.
  • Incorporation of external knowledge graphs to augment vocabulary breadth and enhance semantic precision in disaster-related narratives.
  • Comprehensive disaster semantic alignment and caption informativeness evaluation utilizing CLIPScore and InfoMetIC.
Our VLCE framework produces concise and useful descriptions of disaster imagery, aiding disaster responders, researchers, environmental scientists, and GIS specialists in swift post-disaster evaluation and analysis. The technology will aid NGOs, local communities, and visually impaired individuals by offering accessible summaries that improve communication, decision-making, and situational awareness in emergency situations.
2 Related Works
Researchers have explored numerous studies in this field. Among them, some few works closely related to our research objectives are discussed below:
Author of Zhai et al.[33] proposes a YOLOv8 model for the detection of UAVs using SPD-Conv, and GAM attention with precision, recall, and mAP by 11.9%, 15.2%, and 9%. Similarly, the author of Wang et al.[34] highlights a UAV aerial image on YOLOv8 model includes with WIoU v3 for better localisation, BiFormer attention for better feature attention, and a FasterNet block with new detection with a 7.7% accuracy. Another author[35] proposes six tuned YOLOv8 models for specific ranges of object sizes and performance detection by (mAP-50).
A new study by Abbas et al.[14] investigates CNN-based image caption generation in disaster areas and achieves a BLEU-1 score of 0.8731 and a CIDEr score of 5.0908. Other authors Chun et al.[15] suggested deep learning-based models and obtained an accuracy of 92.9% for damage descriptions. Computer vision plays a vital role in the detection of disasters[9]. The authors in [11] propose a two-step CNN approach that identifies buildings on the xBD dataset scoring with 0.66 F1. Gupta et al.[24] introduced a dataset named xBD for damage assessment. While another author also[7] proposes a UAV-based VQA system for post-disaster damage assessment with a test accuracy of 0.58. However, some researchers use multimodal techniques and combine image and text data to improve disaster image classification with captioning[25].
Recently, image captioning systems have been improving by adding external knowledge[19] help to solve domain-specific problems[6], and refining the model to make context clear. In a different study by Cornia et al.[21] explores transformer architecture for better context using the MS COCO dataset with CIDEr on 132.7. As part of context understanding tasks,Liu et al.[26]offer a GPT-like model for visual instruction tuning, integrating vision and language transformers. This review[27] examines DL approaches for image captioning in diverse applications with a CIDEr score of 15.04%, However, this author[28] provides a pre-trained CNN model for object feature analysis on MS COCO and Flickr datasets. Alternatively, this paper[36] presents InfoMetIC, a reference-free image caption quality metric that provides text accuracy, quality, and vision recall, along with fine-grained feedback of incorrect words. Same as another paper[37] also introduces CLIPScore, a reference-free image caption quality metric using the pretrained CLIP model. LLaVA[26] and Qwen-VL[43] are the vision-language models to generate accurate, context-aware captions and multimodal reasoning.
On the other hand, researchers are now using methods like multi-modal knowledge graphs and transformer models for better context understanding[16, 17, 18]. Despite these efforts, Yao et al.[23] discussed image captioning by leveraging knowledge graphs with MS COCO and Visual Genome datasets. This study discusses[20] commonsense to visual reasoning, challenges, and understanding for better context. Within the object identification, YOLOv8 model[22] helps to improve accuracy, speed, and versatility. In a paper, Zhao et al.[29] suggested a method on multi-modal knowledge graphs that enhances captioning for graph attention networks. Another research also demonstrated[30] better results on CIDEr-D and ROUGE-L to enrich knowledge and more contextually accurate captions. A KG-guided attention mechanism is introduced[31] by adding knowledge triples for semantic and visual balancing in CNN and LSTM models. A transformer-based decoding and graph aggregation model is introduced on MS COCO and Flickr30k datasets[32]. Embeddings are also very important in image captioning. This paper[39] compares different GloVe, BERT, and TaCL for generating captions whereas another paper[40] compares word embeddings for similarity tasks. Furthermore, another paper[41] examines ConceptNet embeddings as background knowledge enhancement for commonsense reasoning.
3 Datasets
We have used the xBD dataset [24] and the RescueNet dataset [42] for our experimental evaluation. The xBD dataset contains 12,738 images, whereas we utilized 6,369 post-disaster images with building damage annotations from five damage categories for our experiment. For our experimentation purpose, we merged two labels, "major-damage" and "minor-damage," into one label, the "damaged" class. This preprocessing phase created three labels: "no-damage," "damaged," and "destroyed." We split our dataset into an 80:20 ratio, where 80% is used for training (5,095 images) and 20% for testing (1,274 images).
[…]
For both datasets, we generate captions using two state-of-the-art vision-language models:
LLaVA-Generated Captions: where Captions are created by the LLaVA model, and, QwenVL-Generated Captions: where captions are generated by the QwenVL model.
4 Proposed Methodology
This section outlines the methodology of the proposed Vision Language Caption Enhancer (VLCE) framework for captioning disaster imagery. We first conduct YOLO object detection to extract identified objects for prompting vision-language models (LLaVA and QwenVL) in the generation of captions. Initially, we outline the two baseline image description generation processes utilizing LLaVA and QwenVL, followed by a presentation of the two model architectures employed in VLCE. A CNN-LSTM model employing a ResNet50 backbone and a Vision Transformer (ViT)-based model are presented.
4.1 Scene-Aware Captioning via Vision-Language Models
To improve the quality of caption creation, we incorporate two advanced vision-language models, LLaVA and QwenVL, to establish our baseline pipeline. Both algorithms utilize object-level priors derived from YOLOv8, an efficient and lightweight object detection framework. For each disaster picture $I$, YOLOv8 generates an annotation set $\mathcal{A}=\{(b_{i},\ell_{i})\}_{i=1}^{K}$, where $b_{i}\in\mathbb{R}^{4}$ represents the coordinates of the bounding box and $\ell_{i}\in\mathcal{L}$ signifies the class label of the $i$-th identified object. The annotated object information is contextually integrated into a fixed textual prompt $P(\mathcal{A})$, which delineates the scene (e.g., “destroyed buildings, flooded road”) and directs the vision-language captioning models.
4.1.1 Context-Aware Captioning with LLaVA
We utilize LLaVA-7B v1.5, which integrates a CLIP ViTL/14 visual encoder with a causal LLaMA decoder within an instruction-following architecture. For an input image $I$, the ViT encoder and its corresponding processor $\phi$ generate a sequence of patch embeddings $V$.
$V=\phi(I)=[v_{1},\ldots,v_{N}]\in\mathbb{R}^{N\times D},$
where $N$ is the number of patches and $D$ the embedding dimension. These visual tokens are prefixed and suffixed with special image markers and concatenated with the disaster context prompt $P(\mathcal{A})$ derived from YOLOv8 annotations into a unified token sequence.
$s=[\texttt{[IM\_START]},\,V,\,\texttt{[IM\_END]},\,P(\mathcal{A})].$
This sequence, $s$ is fed into the LLaMA decoder, which auto-regressively generates the caption $\hat{y}$ by maximizing.
$\hat{y}=\arg\max_{y}P(y\mid s)$
We set the low-temperature decoding(T=0.2) and a maximum length of 4000 tokens.
4.1.2 Context-Aware Captioning with QwenVL
We incorporate the QwenVL-7B Instruct model, a vision-language transformer that tightly integrates a visual encoder with a multilingual LLM, for disaster caption generation. Given an input satellite image $I$, the QwenVL vision encoder processes the image through its built-in processor to produce visual representations $V$.
$V=\text{QwenVL}_{\text{vision}}(I)\in\mathbb{R}^{H\times W\times D},$
where $H$, $W$, and $D$ represent the spatial dimensions and feature depth of the visual encoding. The disaster context is formulated as a structured prompt $P(\mathcal{A},e)$ that incorporates both the predicted annotations $\mathcal{A}$ from YOLOv8 and the event type $e$ extracted from the image metadata:
[…]
The visual and textual inputs are structured as a multi-modal sequence:
$s=\left[\{\text{role: user},\text{content: }[V,P(\mathcal{A},e)]\}\right]$
The unified sequence $s$ is processed by QwenVL’s language model, which generates the detailed disaster caption $\hat{y}$ through autoregressive decoding:
$\hat{y}=\arg\max_{y}\,P(y\mid V,P(\mathcal{A},e))$
We utilized greedy decoding with a maximum generation length of 128 tokens to ensure focused and concise disaster descriptions.
4.2 VLCE Architecture Captioning with Contextual Enrichment
This study proposes generating captions from satellite pictures utilizing the VLMs (LLaVA and QwenVL) model. These captions are augmented by our suggested VLCE architecture (utilizing two methodologies: Transformer (ViT) and LSTM (ResNet50-Eurosa)), which yields more precise, contextually aware, and enriched descriptions for improved catastrophe evaluation. VLCE constitutes our comprehensive model architecture. It incorporates object detection, keyword extraction, knowledge graph (KG) integration, and cross-modal attention to produce contextually relevant image captions from raw photos, resulting in enhanced descriptions. The complete end-to-end design of our innovative solution is depicted below as an algorithm 1.
Algorithm 1 Knowledge Graph-Enhanced VLM Caption Refinement for Disaster Images
1:Disaster image dataset $\mathcal{D}=\{I_{1},\dots,I_{n}\}$
2:Enhanced captions $\mathcal{C}_{\text{Enhanced}}$ with improved semantic accuracy
3:
4:Stage 1: Initial Caption Generation
5:for each image $I_{i}\in\mathcal{D}$ do
6:  Detect objects: $B_{i}\leftarrow\text{YOLO}(I_{i})$
7:  Extract labels: $O_{i}\leftarrow\text{ExtractObjects}(B_{i})$
8:  Construct prompt: $P_{i}\leftarrow\text{ConstructPrompt}(I_{i},O_{i})$
9:  Generate caption: $C_{i}^{\text{VLM}}\leftarrow\text{VLM}(I_{i},P_{i})$
10:end for
11:
12:Stage 2: Caption Preprocessing and Knowledge Preparation
13:Clean captions: $\mathcal{C}_{\text{Clean}}\leftarrow\text{Preprocess}(\{C_{i}^{\text{VLM}}\})$
14:Build vocabulary: $\mathcal{V}\leftarrow\text{BuildVocabulary}(\mathcal{C}_{\text{Clean}})$
15:Create image-caption dictionary: $\mathcal{M}\leftarrow\{(I_{i},c_{i})\}$
16:
17:Stage 3: Feature Extraction and Knowledge Graph Construction
18:Extract visual features: $\mathcal{F}_{\text{Img}}\leftarrow E_{\text{Visual}}(\mathcal{D})$
19:Build KG and embeddings: $\mathcal{G}\leftarrow\text{ConstructKG}(\mathcal{V},\mathcal{C}_{\text{Clean}})$
20:Learn embeddings: $\mathcal{E}_{\text{KG}}\leftarrow\text{KGEmbedding}(\mathcal{G})$
21:Enrich vocabulary: $\mathcal{V}_{\text{Enriched}}\leftarrow\mathcal{V}\cup\text{ExtractRelations}(\mathcal{E}_{\text{KG}})$
22:
23:Stage 4: Multimodal Model Training
24:Initialize multimodal network: $\mathcal{N}_{\theta}=\text{MultimodalNN}(\mathcal{F}_{\text{Img}},\mathcal{V}_{\text{Enriched}},\mathcal{E}_{\text{KG}})$
25:Optimize: $\theta\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}$ over dataset $\mathcal{M}$ with cross-entropy and KG loss
26:
27:Stage 5: Enhanced Caption Generation
28:for each new image $I_{i}\in\mathcal{D}$ do
29:  $C_{i}^{\text{Enhanced}}\leftarrow\text{GenerateCaption}(I_{i},\mathcal{N}_{\theta},\mathcal{V}_{\text{Enriched}},\mathcal{E}_{\text{KG}})$
30:  $\mathcal{C}_{\text{Enhanced}}\leftarrow\mathcal{C}_{\text{Enhanced}}\cup\{C_{i}^{\text{Enhanced}}\}$
31:end for
32:return $\mathcal{C}_{\text{Enhanced}}$
Our system utilizes a structured data architecture in which disaster image datasets and their corresponding caption descriptions are imported from CSV files. Image filenames function as unique identifiers within a dictionary, with associated captions as values. The captioning system preprocesses input data by normalizing text (converting to lowercase), removing punctuation, eliminating single-character artifacts, and retaining just alphabetic tokens. This preprocessing guarantees that the vocabulary comprises significant terms and aids in the generation of embeddings. The comprehensive baseline design is depicted in Figure 2.
Figure 2: Contextual Captioning with Knowledge Graphs
Our knowledge-guided captioning system’s general workflow is shown in Figure 2. YOLOv8 is used for object detection in order to locate items associated with disasters, such as vehicles, debris, and wrecked structures. Through context-aware prompts that describe scene conditions and detected items, the detected objects are integrated into vision-language model prompts to provide context-specific and enriched captions.
4.2.1 Keyword Extraction
RAKE (Rapid Automatic Keyword Extraction) is utilized for keyword extraction by identifying multi-word patterns through frequency analysis and co-occurrence patterns, while excluding short or non-alphabetic words. This process identifies information-dense terms from captions, emphasizing disaster-specific concepts such as "debris field," "structural damage," or "emergency response," which possess greater semantic relevance than isolated terms. In vocabulary construction, we establish word-to-index mappings and rank disaster-related keywords (e.g., "sustained minor," "caused significant," "affected area") to develop a refined vocabulary set for model training. The extracted keywords serve as the basis for knowledge graph integration.
4.2.2 Related Word Formation Using Wordnet and Conceptnet
We utilize related word construction to enhance vocabulary with semantically associated terms, thereby augmenting the model’s comprehension of catastrophic scenarios through the merging of WordNet and ConceptNet.
Keywords from captions are initially enhanced using WordNet for synonym extraction, yielding alternative terms with analogous meanings. Semantic categories classify nouns based on significance (e.g., artifacts, locations, objects), while relationship mapping establishes linkages across disaster-related notions. This lexical enhancement allows the model to acquire synonyms, hypernyms, and hyponyms of extracted terms, improving expressiveness and adaptability.
To further enrichment, we utilize ConceptNet to define node-edge links among keywords and develop semantic connections inside embeddings. For instance, "hurricane" is associated with "wind," "flooding," and "evacuation." We utilize API queries to obtain pertinent terms and connection types, employing relation filtering that omits synonyms to concentrate on conceptual ties. The comprehensive vocabulary includes 1,566 semantically associated terms. This hybrid methodology employing WordNet and ConceptNet enhances vocabulary breadth and augments semantic comprehension.
4.2.3 Knowledge Informed Embedding
We incorporate knowledge-informed textual embeddings into our model to collect contextual information. In particular, to improve relational understanding, we integrate WordNet with 300-dimensional ConceptNet embeddings that maintain high-level semantic relationships. This combination enables the architecture to enhance generated captions with domain-relevant information while preserving semantic consistency across various entities.
4.2.4 Image Feature Extraction
We utilize two distinct architectures for visual feature extraction. The first model is a Transformer-based architecture employed for drone imaging, incorporating a Vision Transformer (ViT) backbone derived from the Remote-Sensing-UAV-Image-Classification paradigm [^1] https://huggingface.co/SeyedAli/Remote-Sensing-UAV-image-classification. The second model is built on CNN-LSTM architecture, utilizing satellite imagery and employing a ResNet50 backbone derived from the ResNet50-EuroSAT model [^2] https://huggingface.co/cm93/resnet50-eurosat. We chose the CNN-LSTM architecture for satellite images due to its pre-training on satellite imagery datasets. The ViT model produces 768-dimensional feature vectors utilizing patch-based attention to encapsulate both global and local contexts. The CNN architecture employs pre-trained ImageNet weights, omitting the classifier, to derive intermediate feature representations prior to the final classification layer. Image preprocessing encompasses scaling, array transformation, and the implementation of a CNN-specific preprocessing pipeline. The selection of an appropriate neural network for transfer learning is essential. ResNet50 (ResNet50-EuroSAT) pulls features from unprocessed pictures, whereas pre-trained embeddings like ConceptNet manage textual features.
4.3 Sequential Modeling
Our framework amalgamates two architectures—Transformer and LSTM—to encapsulate sequential dependencies. Image features, represented as $\mathbf{F}_{\text{Image}}^{T,L}$, are analyzed utilizing a Transformer architecture for drone photos and an LSTM architecture for satellite images. Text features are integrated through knowledge graphs, represented as $\mathbf{F}_{\text{Text}}^{\text{KG}}$ (ConceptNet and WordNet), to enhance semantic information. Visual and textual features are processed using encoder-decoder layers and integrated using $f_{\text{Fusion}}$: an addition layer for the LSTM architecture or a projection layer for the Transformer architecture. The aligned features create the knowledge-aware multimodal feature space, $\mathbf{F}_{\text{Multimodal Feature Space}}$, utilized for caption synthesis. This fusion is mathematically expressed as:
$\mathbf{F}_{\text{Multimodal Feature Space}}=f_{\text{fusion}}\big(\mathbf{F}_{\text{image}}^{\{T,L\}},\mathbf{F}_{\text{text}}^{\text{KG}}\big)$
The resulting knowledge-aware embeddings, enriched vocabulary with WordNet and ConceptNet for better relational information, are combined with visual features extracted from backbones such as the Vision Transformer (ViT) or LSTM (ResNet).
4.3.1 Transformer Architecture
To produce contextually accurate image captions, we develop a hierarchical cross-modal transformer architecture that makes use of semantically enriched textual embeddings and multi-scale visual representations. Our method tackles two major issues in picture captioning: (i) the semantic discrepancy between linguistic representations and visual features, and (ii) the requirement for multi-granularity visual comprehension to capture both fine-grained object details and broad scene context.
Architecture Overview
Our architecture employs a dual-pathway design that includes a hierarchical visual encoder and a cross-modal transformer decoder, as depicted in Figure 3. Rather than employing single-scale visual elements typical of conventional approaches, we explicitly utilize model spatial hierarchies to collect semantic information across many levels.
Hierarchical Visual Encoding
Input image feature vector $\mathbf{I}\in\mathbb{R}^{d_{\text{img}}}$ extracted from a vision transformer backbone where we construct a multi-scale visual representation through parallel projection pathways:
Global Semantic Context: We develop a comprehensive scene representation through dense projection:
$\mathbf{f}_{\text{global}}=\text{ReLU}(\mathbf{W}_{g}\mathbf{I}+\mathbf{b}_{g})\in\mathbb{R}^{d_{\text{model}}}$ (1)
Regional Spatial Context: We project and restructure features into regional patches in order to capture intermediate spatial relationships:
$\mathbf{f}_{\text{regional}}=\text{Dense}_{d_{\text{model}}}(\text{Reshape}_{4\times d_{\text{model}}/4}(\text{ReLU}(\mathbf{W}_{r}\mathbf{I}+\mathbf{b}_{r})))\in\mathbb{R}^{4\times d_{\text{model}}}$ (2)
Local Detail Context: To obtain detailed object and texture information, we develop local feature representations:
$\mathbf{f}_{\text{local}}=\text{Dense}_{d_{\text{model}}}(\text{Reshape}_{12\times d_{\text{model}}/12}(\text{ReLU}(\mathbf{W}_{l}\mathbf{I}+\mathbf{b}_{l})))\in\mathbb{R}^{12\times d_{\text{model}}}$ (3)
The global feature is restructured and combined with regional and local features.
$\mathbf{F}_{v}=\text{LayerNorm}([\text{Reshape}_{1\times d_{\text{model}}}(\mathbf{f}_{\text{global}});\mathbf{f}_{\text{regional}};\mathbf{f}_{\text{local}}])\in\mathbb{R}^{17\times d_{\text{model}}}$ (4)
This hierarchical design enables the model to attend to visual information at appropriate granularities during caption generation, from scene-level understanding to object-specific details.
Cross-Modal Transformer Decoder
Our transformer decoder architecture allows visual-textual alignment while maintaining linguistic structure by combining positional information with pre-trained semantic embeddings.
Semantic Embedding Integration: Input tokens are embedded using a frozen pre-trained embedding matrix $\mathbf{E}\in\mathbb{R}^{|V|\times d_{\text{emb}}}$ to preserve semantic relationships which is learned from large-scale text corpora:
$\mathbf{H}^{(0)}=\mathbf{E}[\mathbf{w}_{1:T}]+\text{PE}(\mathbf{w}_{1:T})$ (5)
where $\text{PE}(\cdot)$ denotes sinusoidal positional encodings that maintain temporal sequence structure.
Multi-Layer Attention Mechanism: The decoder employs $L$ transformer layers for each containing:
  1. 1.
    Masked Self-Attention with causal masking for autoregressive generation:
    $\mathbf{A}^{(l)}_{\text{self}}=\text{MultiHead}(\mathbf{Q}^{(l)},\mathbf{K}^{(l)},\mathbf{V}^{(l)},\mathbf{M}_{\text{causal}})$ (6)
  2. 2.
    Cross-Modal Attention that aligns textual queries with hierarchical visual features:
    $\mathbf{A}^{(l)}_{\text{cross}}=\text{MultiHead}(\mathbf{H}^{(l)},\mathbf{F}_{v},\mathbf{F}_{v})$ (7)
  3. 3.
    Position-wise Feed-Forward Networks with residual connections:
    $\mathbf{H}^{(l+1)}=\text{LayerNorm}(\mathbf{A}^{(l)}_{\text{cross}}+\text{FFN}(\mathbf{A}^{(l)}_{\text{cross}}))$ (8)
This design enables the model to simultaneously capture intra-sequence dependencies and visual-textual correspondences.
Multi-Modal Fusion and Output Generation
We utilized a fusion strategy that integrates global visual semantics with local textual representations to improve the final predictions with global visual context:
$\mathbf{c}_{\text{visual}}$ $=\operatorname{GlobalAvgPool}(\mathbf{F}_{v}),$ (9)
$\hat{\mathbf{y}}_{t}$ $=\operatorname{Dense}_{V}\Big(\operatorname{LayerNorm}\big([\mathbf{h}_{t}^{(L)};\mathbf{c}_{\text{visual}}]\big)\Big),$ (10)
where $\cdot$$\cdot$] denotes concatenation and $\operatorname{Dense}_{V}$ projects to the vocabulary space.
Training Objective and Optimization
We integrated a masked sparse categorical cross-entropy loss that focuses learning on meaningful tokens while ignoring padding:
$\mathcal{L}=-\frac{1}{\sum_{t=1}^{T}m_{t}}\sum_{t=1}^{T}m_{t}\log P(w_{t}|\mathbf{w}_{<t},\mathbf{F}_{v};\theta)$ (11)
where $m_{t}\in\{0,1\}$ is a mask indicating non-padding positions, and $\theta$ represents all model parameters.
Figure 3: Hierarchical cross-modal transformer architecture
Our architecture utilizes with $d_{\text{model}}=d_{\text{emb}}=300$ for dimensional consistency, $L=2$ transformer layers with $h=6$ attention heads each one. We implement a two-phase training strategy: initial exploration with higher learning rates followed by fine-tuning at reduced rates. Dynamic batch generation with prefetching ensures optimal GPU utilization during training.
Our proposed hierarchical cross-modal transformer effectively connects the semantic gap between visual and linguistic modalities while capturing the multi-scale spatial dependencies, enabling robust caption generation for complex visual scenes with enhanced semantic accuracy.
4.3.2 LSTM Architecture
We employ a CNN-LSTM architecture for image captioning that utilizes pre-extracted picture attributes and textual embeddings. The architecture has two main components: a visual feature extractor and a sequential language model.
The visual encoder processes image features $I\in\mathbb{R}^{d_{\text{img}}}$ through a dropout layer and a fully connected layer to yield a 256-dimensional feature vector.
$f_{\text{image}}=\text{Dense}_{256}(\text{Dropout}_{0.5}(I))$
The pipeline of LSTM based architecture is shown in Figure 4.
Figure 4: LSTM architecture with visual and textual branches
Figure 4 depicts the LSTM-based image captioning architecture, wherein the image branch extracts and encodes visual information, while the text branch processes the caption input through embeddings.
In the textual decoder, captions are tokenized and embedded utilizing a pre-trained embedding matrix $E\in\mathbb{R}^{|V|\times d_{\text{emb}}}$, which can be selectively frozen to maintain acquired semantic associations. The embeddings undergo regularization using dropout and are subsequently processed by an LSTM layer containing 256 hidden units.
$h_{\text{text}}=\text{LSTM}_{256}(\text{Dropout}_{0.5}(E[w_{1:T}]))$
The projected image features and LSTM output are integrated through an additive operation, succeeded by a dense layer and culminating in a final softmax layer applied to the vocabulary:
$\hat{y}_{t}=\text{Softmax}(\text{Dense}_{|V|}(\text{Dense}_{256}(f_{\text{image}}+h_{\text{text}})))$
The training process utilizes a specialized data generator that creates tuples consisting of image features, input sequences, and target sequences for every batch, maintaining a batch size of 32. The network employs categorical cross-entropy loss for compilation and utilizes the Adam optimizer for enhancement. The architecture incorporates image feature dimensions of $d_{\text{img}}=2048$ and embedding dimensions of $d_{\text{emb}}=300$, with dropout implemented in both the image and text branches. The size of the vocabulary $|V|$ is influenced by the specific dataset and the choices made during preprocessing. This architecture effectively combines pre-trained semantic embeddings with visual features extracted by CNNs to produce precise captions.
4.4 Generating captions
Automatic image captioning produces coherent and contextually relevant natural language descriptions of visual content. This study employs a combination of transformer-based and LSTM-based decoding methodologies. Both approaches utilize encoded image representations as input and generate sequential textual outputs.
Our initial method utilizes a vision transformer encoder to derive high-dimensional image embeddings, succeeded by a transformer decoder for autoregressive text generation. For an image $I$, the ViT encoder generates a sequence of patch embeddings $E=\{e_{1},e_{2},\dots,e_{n}\}$. The decoder produces a caption $C=\{w_{1},w_{2},\dots,w_{T}\}$ by forecasting each word based on the image embeddings and all words generated prior:
$w_{t}=\arg\max_{w\in V}P(w\mid w_{1},\dots,w_{t-1},E)$
where $V$ represents the vocabulary, and $t\in[1,T]$. The model employs post-padding to standardize input sequences to the maximum length. In the process of inference, a START token is employed to initialize the sequence, and the generation proceeds iteratively until either a STOP token is predicted or the maximum sequence length is attained. The architecture leverages the self-attention mechanism of transformers, which adeptly captures long-range dependencies between visual and textual modalities.
Our second approach employs a convolutional neural network (CNN) as an image encoder to derive a global visual feature vector. This feature is integrated into a long short-term memory network to generate sequential word predictions. In a formal manner:
$w_{t}=\arg\max_{w\in V}P(w\mid w_{1},\dots,w_{t-1},f_{I})$
where $f_{I}$ denotes the encoded visual feature derived from the CNN. The LSTM decoder produces one word sequentially, preserving temporal dependencies through its internal hidden state. Sequence padding guarantees uniform input length, and generation concludes upon the detection of the STOP token.
Both architectures utilize wordtoidx and idxtoword mappings for vocabulary management. The special tokens START and STOP indicate the boundaries of a series. The transformer necessitates indexing the current prediction position because it generates full-sequence outputs, while LSTM produces single-step predictions. The transformer architecture employs sparse categorical cross-entropy loss, whereas the LSTM architecture utilizes categorical cross-entropy loss. Captions are produced incrementally by adding predicted words until the STOP token is encountered.
5 Experimentation
5.1 Dataset Preprocessing
The data preprocessing pipeline processes post-disaster imagery obtained from satellite and UAV sources. Captions are produced utilizing the LLaVA and QwenVL models. Data is organized as image-caption dictionaries to facilitate efficient model training and evaluation.
Image Preprocessing: In Vision Transformer experiments, photos are standardized, resized to $224\times 224$ pixels, transformed to RGB format, and normalized to the $[0,1]$ range. Feature extraction employs global average pooling on the outputs of the last transformer layer, resulting in 768-dimensional feature vectors. In CNN-based studies, photos are enlarged to $299\times 299$ pixels with pretrained ResNet models, resulting in 2,048-dimensional feature representations.
Caption Preprocessing: Text normalization encompasses tokenization, conversion to lowercase, removal of punctuation, and the exclusion of non-alphabetic tokens that are shorter than two characters. Artifacts and stop words are removed. Captions utilize boundary tokens startseq and endseq, adhering to a maximum sequence length of 192 tokens.
Vocabulary Construction: The initial vocabulary consists of distinct tokens derived from training captions. In knowledge graph investigations, vocabulary is augmented through keyword extraction utilizing RAKE, succeeded by semantic expansion with ConceptNet and WordNet APIs. Associated terms are obtained while omitting synonyms to save redundancy. Duplicate terms and border tokens are eliminated, while overlap removal mitigates semantic redundancy.
Embedding Integration: ConceptNet Numberbatch 300-dimensional vectors with random initialization for missing terms and special tokens are used in knowledge graph research. Vocabulary alignment is easier with prefix matching. BERT embeddings (768-dimensional vectors) are used in baseline experiments without knowledge graphs. The 80:20 train-test ratio splits both datasets for consistent evaluation across experiments.
5.2 Evaluation Metrics
We utilize two complementing evaluation techniques to assess caption quality: CLIPScore for semantic alignment and InfoMetIC for informativeness assessment.
CLIPScore Evaluation: CLIPScore evaluates semantic alignment between generated captions and images by computing the cosine similarity of CLIP embeddings.
$\text{CLIPScore}=\cos\!\big(\text{CLIP}_{\text{image}}(I),\text{CLIP}_{\text{text}}(C)\big)$
where $I$ denotes the input image and $C$ indicates the generated caption. The assessment procedure is delineated in Algorithm 2.
Algorithm 2 CLIPScore-Based Evaluation of Disaster Image Captions
1:Custom captions $\mathcal{C}_{A}$, Baseline captions $\mathcal{C}_{B}$, Image set $\mathcal{I}$, CLIP model
2:CLIP scores $\mathcal{S}_{A},\mathcal{S}_{B}$ and comparison statistics
3:
4:Stage 1: Initialization
5:Load CLIP model and processor
6:Align captions $\mathcal{C}_{A},\mathcal{C}_{B}$ with images $\mathcal{I}$
7:
8:Stage 2: CLIPScore Computation
9:for each image $I_{i}\in\mathcal{I}$ do
10:  $s_{i,A}\leftarrow\text{CLIPScore}(I_{i},C_{i,A})$
11:  $s_{i,B}\leftarrow\text{CLIPScore}(I_{i},C_{i,B})$
12:  Store $(I_{i},s_{i,A},s_{i,B})$ in $\mathcal{R}$
13:end for
14:
15:Stage 3: Comparative Analysis
16:Count $n_{\text{better}}\leftarrow|\{i:s_{i,A}>s_{i,B}\}|$
17:Compute percentage $p=\dfrac{n_{\text{better}}}{|\mathcal{I}|}\times 100$
18:Output results $\mathcal{R}$ and $p$
InfoMetIC Evaluation: InfoMetIC assesses caption informativeness by combining relevance, informativeness, and precision components:
$\text{InfoMetIC}=\alpha\times\text{Informativeness}+\beta\times\text{Relevance}+\gamma\times\text{Precision}$
where $\alpha$, $\beta$, and $\gamma$ are weighting factors determined during validation. This metric is particularly valuable for disaster image captioning as emergency responders require detailed, relevant damage descriptions. The evaluation process is detailed in Algorithm 3.
Algorithm 3 InfoMetIC-Based Evaluation of Disaster Image Captions
1:Caption CSVs: $\text{CSV}_{\text{Custom}},\text{CSV}_{\text{Baseline}}$; image directory $\mathcal{I}$; CLIP model; corpus frequency dictionary
2:InfoMetIC scores and comparative statistics
3:
4:Stage 1: Load Model and Data
5:Load CLIP model and processor
6:Configure device: GPU if available, else CPU
7:Load $\text{CSV}_{\text{Custom}}$ and $\text{CSV}_{\text{Baseline}}$
8:Filter both CSVs to include only images present in $\mathcal{I}$
9:
10:Stage 2: Compute InfoMetIC Scores
11:for each image–caption pair $(i,c)\in\text{CSV}_{\text{Custom}}\cup\text{CSV}_{\text{Baseline}}$ do
12:  $R(i,c)\leftarrow\mathrm{CLIPScore}(i,c)$ $\triangleright$ Relevance
13:  $I(c)\leftarrow\sum_{w\in c}-\log P(w)$ $\triangleright$ Informativeness from corpus
14:  $S(i,c)\leftarrow R(i,c)\times I(c)$ $\triangleright$ InfoMetIC score
15:  Store $(i,S(i,c))$ in results table
16:end for
17:
18:Stage 3: Merge and Compare
19:Merge results for $\text{CSV}_{\text{Custom}}$ and $\text{CSV}_{\text{Baseline}}$ by image ID
20:Define comparison indicator for each image $i$:
$\text{Better}(i)=\begin{cases}1&\text{if }S_{\text{Custom}}(i)>S_{\text{Baseline}}(i)\\ 0&\text{otherwise}\end{cases}$
21:Compute performance percentage:
$\text{Percentage}_{\text{Better}}=\frac{\sum_{i}\text{Better}(i)}{\text{Total Images}}\times 100$
Complementary Evaluation Framework: InfoMetIC assesses captions using text precision for word accuracy, vision recall for completeness, and a combined score. CLIPScore may give generic but relevant captions high scores (e.g., "a building" for disaster imagery). InfoMetIC penalizes uninformative descriptions with word rarity, identifying semantically accurate and information-rich captions.
The implementation employs CLIP for relevance computation, ensuring consistency with CLIPScore evaluation, and utilizes Brown corpus frequencies for the assessment of informativeness. This method guarantees methodological consistency within our evaluation framework and enhances computational efficiency for large-scale analyses. The dual-metric evaluation facilitates a thorough assessment of semantic alignment and descriptive richness in disaster image captions.
5.3 Embedding Process With KG and Without KG
We integrated two approaches for generating textual embeddings for disaster image captions: one incorporating knowledge graph information and another is without KG integration. In the KG-based approach, we utilized ConceptNet and WordNet embeddings, and in the without-KG approach, we employed BERT embeddings.
5.3.1 With ConceptNet and WordNet embedding
To improve semantic coverage for captioning disaster images, we integrated embeddings from WordNet and ConceptNet for vocabulary expansion of the disaster-related words. Our pipeline extracted relevant keywords and related terms from training captions and embedded them to build a knowledge-informed vocabulary.
For WordNet integration, we used the WordNet lexical database to synonyms extraction for each keyword in the top ten phrases per caption. We only considered only the first word of each keyword phrase. We collected all unique synonyms and create a dictionary, filtering out non-alphabetic terms and duplicates. To prevent semantic redundancy, overlapping substrings were removed, resulting in 1,566 WordNet-based semantic words processed for the overall caption lexicon.
ConceptNet embeddings were integrated using Numberbatch vectors, which offer 300-dimensional dense representations for words and phrases. Each term $t\in V$ was mapped to a vector:
$\mathbf{e}_{t}\in\mathbb{R}^{300}.$
where $\mathbf{e}_{t}$ is the embedding vector of term $t$. Given the full embedding dictionary $D=\{(w,\mathbf{e}_{w})\}$, Prefix matching was used to match vocabulary terms with their corresponding ConceptNet embeddings.
$\mathbf{e}_{t}=\begin{cases}\text{find}(D,t)&\text{if }t\in D,\\[6.0pt] \mathbf{e}_{\text{rand}}\sim\mathcal{U}(-\epsilon,\epsilon)&\text{otherwise}.\end{cases}$
where $\mathbf{e}_{\text{rand}}$ is a randomly initialized vector sampled from a uniform distribution $\mathcal{U}(-\epsilon,\epsilon)$. Special tokens (such as startseq and endseq) and words were initialized using random vectors. The final embedding matrix was constructed as:
$E=\begin{bmatrix}\mathbf{e}_{t_{1}}\\ \mathbf{e}_{t_{2}}\\ \vdots\\ \mathbf{e}_{t_{|V|}}\end{bmatrix}\in\mathbb{R}^{|V|\times 300},$
where $E$ is the embedding matrix, $t_{i}$ is the $i$-th token in the vocabulary, and $|V|=3195$ is the vocabulary size. Both the original vocabulary and terms enhanced by knowledge graphs were included in the final embedding matrix, which contained 3,195 tokens with 300-dimensional representations.
WordNet and ConceptNet-derived terms were combined to create the knowledge graph-enhanced vocabulary. In addition to remove duplicate entries, boundary tokens (startseq, endseq) were added. A maximum of 192 tokens for the caption was incorporated by sequence padding. After this combination,the model was able to more accurately represent the semantic relationships between disaster-related entities, objects, and locations for the expanded vocabulary.
To create training batches dynamically, a custom Keras data generator was used. For autoregressive caption prediction, each batch included input sequences, target sequences, and image features. The generator supported quality filtering, balancing, and multi-dataset integration, and input sequences were padded to the maximum length. This configuration reduced memory overhead and allowed for effective training with a batch size of 32. The knowledge graph-based vocabulary’s embedding matrix was constructed by the ConceptNet Numberbatch embeddings with WordNet synonyms. A vector with 300 dimensions was assigned to every vocabulary token. Random initialization was used for tokens. Both knowledge-enriched and caption-derived semantic representations utilized by the validation of the embedding matrix for vocabulary alignment. Among the two embeddings, WordNet contributes to lexical diversity, while ConceptNet contributes to domain-specific context. The resulting embedding matrix could effectively use both knowledge-enriched and caption-derived semantic representations.
5.3.2 With Bert Embedding
BERT embedding is used for text embedding when the knowledge graph is not used. The BERT embedding technique introduced a contextualized approach, which utilized distilbert-base-uncased model to generate embeddings.
The embedding matrix is constructed by tokenizer that converts each textual token into subword tokens, which are then fed into the pre-trained BERT model to produce hidden state vectors. For each input word $w$, the embedding $e_{w}\in\mathbb{R}^{d}$ is obtained using the hidden state corresponding to the [CLS] token:
$e_{w}=\text{BERT}(w)_{[\text{CLS}]}$ (12)
Alternatively, mean pooling over all token embeddings can be applied, providing a flexible representation for multi-token words. We also used embedding matrix construction in BERT embedding where $V$ denote the vocabulary size, and $d$ the embedding dimension determined by the BERT configuration (768 for distilbert-base-uncased). We construct an embedding matrix $E\in\mathbb{R}^{V\times d}$, where each row corresponds to a word embedding:
$E[i,:]=\begin{cases}e_{w}&\text{if }w\in\text{vocabulary}\\ \mathcal{N}(0,0.1)&\text{otherwise}\end{cases}$ (13)
Special tokens, including startseq and endseq, are assigned randomly initialized vectors to preserve network stability. The matrix is cached for efficient reuse in subsequent experiments. To load the pre-trained models and tokenizers with GPU inference, the transformers library is integrated with the BERT embedding model architecture. Each word is tokenized individually, and embeddings are extracted without gradient computation to reduce memory overhead. Missing or unprocessable tokens are initialized with small random vectors drawn from a Gaussian distribution. The result of the embedding matrix exhibits high coverage over the vocabulary, with the majority of words successfully mapped to pre-trained embeddings. We created the BERT embedding matrix for our vocabulary of size $V=\text{vocab\_size}$, resulting in a matrix of dimensions $V\times 768$. Semantic coherence and numerical stability were analyzed in a subset of embeddings. In addition, semantic similarity in embeddings can be measured with cosine similarity:
$\text{sim}(e_{w_{1}},e_{w_{2}})=\frac{e_{w_{1}}\cdot e_{w_{2}}}{\|e_{w_{1}}\|\,\|e_{w_{2}}\|}$ (14)
where $e_{w_{1}}\cdot e_{w_{2}}$ is the product of the embedding vectors, and $\|e_{w_{1}}\|$, $\|e_{w_{2}}\|$ are their Euclidean norms. This measure was utilized to test clustering of disaster-related words (e.g., “building”, “debris”, “hurricane”), indicating semantically related tokens have higher similarity values. Such structural facts verify semantic coherence of embeddings, which improves downstream caption generation tasks.
6 Results Evaluation
6.1 Overview of Experimental Results
We assessed our knowledge graph-enhanced image captioning framework against prominent Vision Language Models (VLMs) utilizing two disaster response datasets: RescueNET (UAV imagery) and xBD (satellite imagery). The evaluation utilized two complementing metrics—CLIPScore for semantic alignment and InfoMetIC for informativeness—coupled with noun-based object analysis to assess caption comprehensiveness. Experiments systematically examined the effects of knowledge graph integration across architectural configurations (CNN-LSTM and Transformer) and visual backbones (ViT-UAV and ResNet50-EuroSat).
Tables 1 and 2 exhibit consolidated performance metrics, indicating that the incorporation of knowledge graphs substantially modifies model behavior, with variations significantly influenced by architectural and dataset attributes. Although knowledge graph integration consistently enhances informativeness, its effect on semantic alignment is more intricate and contingent on setup.
Table 1: CNN-LSTM Architecture Performance (LLaVA Captions)
Dataset KG Backbone CLIPScore (%) InfoMetIC (%)
Custom LLaVA Custom LLaVA
RescueNET With ViT (UAV) 52.95 47.05 54.51 45.49
Without ViT (UAV) 0.56 99.44 1.22 98.78
xBD With ResNet50 51.10 48.90 66.56 33.44
Without ResNet50 55.34 44.66 66.41 33.59
6.2 Architecture-Specific Performance Analysis
6.2.1 CNN-LSTM Architecture Performance
The CNN-LSTM architecture, utilizing LLaVA baseline captions, demonstrated considerable sensitivity to the incorporation of knowledge graphs (Table 1). Utilizing the ViT-UAV backbone in RescueNET, the integration of KG yielded performance metrics that were equivalent between our own model and the LLaVA baseline (CLIPScore: 52.95% versus 47.05%; InfoMetIC: 54.51% versus 45.49%). The absence of knowledge graphs resulted in a significant decline in performance, with the custom model attaining less than 2% on both criteria, underscoring the architecture’s essential reliance on structured information for UAV imagery interpretation.
The xBD dataset disclosed varying dynamics. Utilizing a ResNet50 backbone and knowledge graph integration, our methodology attained significant informativeness (66.56%) while preserving moderate semantic alignment (51.10%). The elimination of knowledge graphs enhanced CLIPScore performance (55.34% compared to 44.66%), indicating that for satellite images characterized by discrete visual patterns, CNN-LSTM may gain from direct visual-linguistic mapping, circumventing intermediary knowledge frameworks.
Table 2: Transformer Architecture Performance (QwenVL Captions)
Dataset KG Backbone CLIPScore (%) InfoMetIC (%)
Custom QwenVL Custom QwenVL
RescueNET With ViT (UAV) 73.64 26.36 95.33 4.67
Without ViT (UAV) 0.22 99.78 0.08 99.92
xBD With ResNet50 60.60 39.40 69.86 30.14
Without ResNet50 22.14 77.86 18.76 81.24
6.2.2 Transformer Architecture Performance
The Transformer architecture, when assessed against the QwenVL baseline, exhibited improved robustness and notable benefits from the incorporation of knowledge graphs (Table 2). Utilizing the ViT-UAV backbone in RescueNET, KG integration attained exceptional informativeness (95.33%—the greatest among all setups) while preserving moderate semantic alignment (73.64%). This imbalance suggests that knowledge graphs empower Transformers to produce detailed, informative descriptions despite suboptimal visual-semantic alignment.
The xBD setup utilizing the ResNet50 backbone produced the most equitable performance profile. Through KG integration, our model attained a CLIPScore of 60.60% and an InfoMetIC of 69.86%, demonstrating an optimal equilibrium between semantic accuracy and descriptive richness. In the absence of KG help, performance declined markedly (CLIPScore: 22.14%; InfoMetIC: 18.76%), albeit to a lesser extent than RescueNET, indicating that the organized attributes of satellite images offer intrinsic semantic assistance.
6.3 Metric-Specific Analysis
6.3.1 CLIPScore Evaluation: Semantic Alignment
CLIPScore study demonstrated dataset-specific patterns in the quality of semantic alignment. In UAV imaging (RescueNET), the inclusion of knowledge graphs was essential for semantic consistency. Distribution histograms (Figures 511) demonstrate that KG-enhanced models produce score distributions centered around elevated values, signifying consistent semantic alignment across disaster scenarios. Architectural comparison unveiled divergent response patterns: CNN-LSTM displayed binary performance—either effectively competing with knowledge graphs or entirely failing in their absence—whereas the Transformer architecture exhibited graceful degradation, preserving baseline functionality even without knowledge graph support. This resilience indicates that attention mechanisms offer implicit semantic organizing that partially offsets the absence of explicit knowledge. Satellite imagery (xBD) exhibited more uniform CLIPScore patterns across setups. The ResNet50 backbone consistently attained a performance range of 48–60% in bespoke models, irrespective of the presence of a knowledge graph, while the score distributions exhibited significant variation. The inclusion of KG resulted in bimodal score distributions, signifying efficient separation between high-confidence and uncertain predictions, which is advantageous for operational deployment.
Table 3: Comparison of relevant object counts from captions across various model configurations
Exp. Configuration Custom Model LLaVA QwenVL Best Model
ViT_UAV_RescueNet Dataset
1 With Knowledge Graph 272 185 201 Custom
2 Without Knowledge Graph 272 178 195 Custom
ResNet-EuroSAT (xBD Dataset)
3 With Knowledge Graph 640 445 463 Custom
4 Without Knowledge Graph 640 431 448 Custom
Performance Summary
Average Improvement over LLaVA +45.2%
Average Improvement over QwenVL +38.7%
Experiments Won 4/4 0/4 0/4
(a) Histogram of CLIPScore with KG
(b) Comparison of CLIPScore with KG
Figure 5: CLIPScore evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with LLaVA and knowledge graph: (a) score distribution, (b) model-wise comparison.
(a) Histogram of CLIPScore without KG
(b) Comparison of CLIPScore without KG
Figure 6: CLIPScore evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with LLaVA without knowledge graph: (a) score distribution, (b) model-wise comparison.
(a) Histogram of CLIPScore with KG
(b) Comparison of CLIPScore with KG
Figure 7: CLIPScore evaluation for ResNet-EuroSAT xBD with LLaVA and knowledge graph: (a) score distribution, (b) model-wise comparison.
(a) Histogram of CLIPScore without KG
(b) Comparison of CLIPScore without KG
Figure 8: CLIPScore evaluation for ResNet-EuroSAT xBD with LLaVA without knowledge graph: (a) score distribution, (b) model-wise comparison.
(a) Histogram of CLIPScore with KG
(b) Comparison of CLIPScore with KG
Figure 9: CLIPScore evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with QwenVL and knowledge graph: (a) score distribution, (b) model-wise comparison.
(a) Histogram of CLIPScore without KG
(b) Comparison of CLIPScore without KG
Figure 10: CLIPScore evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with QwenVL without knowledge graph: (a) score distribution, (b) model-wise comparison.
(a) CLIPScore barchart (with KG)
(b) CLIPScore barchart (without KG)
Figure 11: CLIPScore evaluation for ResNet-EuroSAT backbone, xBD dataset with QwenVL VLM: (a) with knowledge graph, (b) without knowledge graph.
6.3.2 InfoMetIC Evaluation: Caption Informativeness
InfoMetIC scores consistently exhibited the effectiveness of knowledge graph integration in generating informative captions. All KG-supported setups attained InfoMetIC scores surpassing 54%, with Transformer-RescueNET achieving 95.33%, indicating near-complete informational dominance over baseline models.
Analysis of score distribution (Figures 1218) indicates that the integration of knowledge graphs enhances average performance and fundamentally alters the density profiles of caption information. KG-enhanced models consistently provide information-dense captions, as seen by right-skewed distributions and diminished variance relative to baseline models.
The complementing relationship between CLIPScore and InfoMetIC measures is notably important. Configurations exhibiting moderate CLIPScore but elevated InfoMetIC, such as Transformer-RescueNET with KG, suggest that knowledge graphs empower models to generate intricate, contextually rich descriptions that surpass mere visual content—potentially integrating domain-specific insights regarding disaster scenarios, damage typologies, and emergency response considerations.
(a) InfoMetIC comparison with KG
(b) Histogram of InfoMetIC with KG
Figure 12: InfoMetIC evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with LLaVA and knowledge graph: (a) model-wise comparison, (b) score distribution.
(a) InfoMetIC comparison without KG
(b) Histogram of InfoMetIC without KG
Figure 13: InfoMetIC evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with LLaVA without knowledge graph: (a) model-wise comparison, (b) score distribution.
(a) InfoMetIC comparison with KG
(b) Histogram of InfoMetIC with KG
Figure 14: InfoMetIC evaluation for ResNet-EuroSAT xBD with LLaVA and knowledge graph: (a) model-wise comparison, (b) score distribution.
(a) InfoMetIC comparison without KG
(b) Histogram of InfoMetIC without KG
Figure 15: InfoMetIC evaluation for ResNet-EuroSAT xBD with LLaVA without knowledge graph: (a) model-wise comparison, (b) score distribution.
(a) InfoMetIC comparison with KG
(b) Histogram of InfoMetIC with KG
Figure 16: InfoMetIC evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with QwenVL and knowledge graph: (a) model-wise comparison, (b) score distribution.
(a) InfoMetIC comparison without KG
(b) Histogram of InfoMetIC without KG
Figure 17: InfoMetIC evaluation for “Remote-Sensing-UAV-image-classification” RescueNet with QwenVL without knowledge graph: (a) model-wise comparison, (b) score distribution.
(a) InfoMetIC with KG
(b) InfoMetIC without KG
Figure 18: InfoMetIC evaluation for ResNet-EuroSAT xBD with QwenVL: (a) with knowledge graph, (b) without knowledge graph.
6.3.3 Noun-Based Object Analysis
The noun analysis (Table 3) offers substantial evidence for improved object detection and description abilities via knowledge augmentation. Our custom model detected 38.7–45.2% more pertinent objects than baseline models across all setups. The uniformity across datasets and architectures illustrates that the incorporation of knowledge graphs markedly enhances models’ ability to recognize and express visual elements.
Analysis revealed 272 unique objects in UAV imagery and 640 in satellite imaging, with our proprietary model attaining full coverage, whereas baseline models achieved just 65–70% coverage. This thorough object recognition is essential for disaster response applications, as missed details may signify vital infrastructure damage, trapped humans, or dangerous situations.
6.4 Key Findings and Implications
Our thorough assessment uncovers three essential conclusions for knowledge-enhanced disaster image captioning:
Architecture-Knowledge Synergy: The correlation between model architecture and knowledge graph integration is non-linear and contingent upon the dataset. Although Transformers provide greater overall performance, CNN-LSTM systems attain competitive results through efficient knowledge integration, especially for satellite imagery characterized by regular spatial patterns.
Complementary Metric Perspectives: CLIPScore and InfoMetIC assess several characteristics of caption quality. The combination of high InfoMetIC and modest CLIPScore suggests that knowledge-enhanced descriptions exceed visual content, potentially providing superior utility for emergency response compared to solely descriptive captions.
Operational Considerations: Marked performance discrepancies among configurations highlight the necessity of meticulous system design for deployment. The Transformer architecture integrated with knowledge graphs on the xBD dataset constitutes the most resilient configuration, demonstrating superior performance across all metrics and exhibiting smooth degradation in the event of component failure.
The findings indicate that effective disaster image captioning necessitates advanced vision-language models, the integration of domain expertise, suitable architectural selections, and thorough evaluation methods. The consistent superiority of knowledge-enhanced models across various metrics supports our hypothesis that structured knowledge integration is crucial for producing captions that satisfy the complex operational demands of disaster response.
7 Case Studies: Impact of Knowledge Graph on Caption Quality
We experimented with several model configurations and datasets to assess the efficacy of Knowledge Graph (KG) integration. Our investigation evaluates caption quality between KG-enhanced and baseline models utilizing two key datasets: RescueNet and XBD. Several model designs, such as ViT, ResNet-EuroSAT, and other caption generation models (LLaVA and QwenVL), are included in the tests. This section illustrates the distinction between figure captions based on the utilization of KG.
Figure 19: “Remote-Sensing-UAV-image-classification” RescueNet With KG-Enhanced Model LLaVA Output
The KG-enhanced caption delivers coherent and contextually precise descriptions. It accurately identifies flooding as the principal type of damage, observes the existence of submerged debris, and deduces recovery operations from the visible vehicles. The narrative maintains a consistent flow without redundancy or fictitious elements.
Image Caption [19] Caption: the image depicts the aftermath of hurricane michael which caused significant damage to the infrastructure and environment in the affected area the most noticeable feature is the flooded area which has submerged the debris from damaged structures can be seen scattered throughout the scene with some debris piled up in the water the presence of multiple vehicles including cars and trucks indicates that the area is being cleared and assessed for recovery efforts the overall scene conveys sense of destruction and recovery efforts needed to restore the affected area to its previous state.
Conversely, Figure 20 illustrates the baseline model output devoid of KG enhancement, indicating substantial quality deterioration.
Figure 20: “Remote-Sensing-UAV-image-classification” Without KG RescueNet LLaVA Output
Image Caption [20] Caption: the the image shows an aerial view of a residential severely impacted by a natural disaster fallen trees lie across rooftops, stone and streets directly visible structural damage can be seen while the surrounding area is littered with broken branches and debris the stark contrast between intact rooftops and widespreadvegetation damage impacted by a natural disaster fallen trees lie rooftops, stone
The baseline caption demonstrates various failure modes: word duplication (“the the”), nonsensical terms (“stone”) presented without context, (“widespreadvegetation” as a single word), and self-repetition with entire phrases repeating verbatim. These errors illustrate how the lack of organized knowledge results in both syntactic and semantic failures.
We expanded our assessment to the XBD dataset utilizing the ResNet-EuroSAT architecture. Figure 21 illustrates the results enhanced by knowledge graphs.
Figure 21: ResNetEurosat XBD Dataset With KG-Enhanced Model LLaVA Output
Image Caption [21] Caption: the image shows satellite view of forest with clear path through the trees the path is surrounded by trees and there are several trees in the area the pathway appears to be dirt road and it is surrounded by trees the image also shows few buildings in the area which seem to have escaped the aftermath of the disaster the image captures the aftermath of disaster with undamaged areas or buildings still standing amidst the destruction
The KG-enhanced model effectively detects mixed damage scenarios, identifying both forested regions and buildings, while critically differentiating between damaged and undamaged structures. This detailed evaluation is crucial for disaster response planning.
The baseline model devoid of knowledge graph (Figure 22) produces significantly divergent—and concerning—results.
Figure 22: ResNetEurosat XBD Dataset Without KG-Enhanced Model LLaVA Output
Image Caption [22] Caption: the image shows satellite view of field with large crack in the middle likely caused by an earthquake the crack extends from the left side of the image to the right dividing the field into two distinct sections additionally there are two people in the field one near the center and the other towards the right side the presence of these holes suggests that the area has been affected by the disaster andthe animals deaths may have been temporarily closed due to the damage
The baseline caption includes fabricated elements (individuals in the field, animal fatalities) and semantically incoherent phrases (“animal deaths may have been temporarily closed”). These fabrications exemplify a significant failure mode in which the model produces credible yet entirely inaccurate information—a particularly perilous consequence for disaster assessment applications.
To evaluate the applicability of KG benefits across various caption generation architectures, we examined the QwenVL model. Figure 23 illustrates the performance of KG-enhanced QwenVL on RescueNet.
Figure 23: “Remote-Sensing-UAV-image-classification” With KG RescueNet QwenVL Output
Image Caption [23] Caption: the satellite image depicts an area that has been significantly impacted by hurricane the image shows dense concentration of trees that appear to be damaged or fallen indicating the powerful force of the storm the trees are scattered throughout the area with some lying on the ground and others leaning precariously the extent of the destruction is widespread affecting both the natural landscape and built environments the fallen trees and debris suggest that the hurricane had strong impact on the vegetation potentially causing widespread damage to the local ecosystem the presence of debris and fallen trees can obstruct roads and paths making it difficult for emergency services and residents
The KG-enhanced QwenVL caption exhibits advanced environmental comprehension, accurately recognizing vegetation damage patterns, evaluating ecosystem effects, and deriving pertinent conclusions regarding emergency response implications. The description upholds factual accuracy while offering actionable insights.
Figure 24 demonstrates that the baseline QwenVL output displays analogous degradation patterns to those identified in LLaVA.
Figure 24: “Remote-Sensing-UAV-image-classification” Without KG RescueNet QwenVL Output
Image Caption [24] Caption: the satellite photo depicts an area that has been significantly impacted by a food area and causing harm to five people of the destruction is evident from the scattered debris and damaged structures here is detailed description of the visible damage and its implications damaged structures there are two large piles of debris scattered across the area these piles consist of various materials such as trees, wood metal and possibly concrete indicating that multiple buildings have been severely damaged or destroyed the debris
The baseline output exhibits significant semantic inaccuracies, including the illogical phrase "food area and causing harm to five people" and fragmented sentences. The fabrication of precise casualty figures is especially alarming, as such misinformation could lead to severe repercussions in genuine disaster response situations. The concluding assessment analyzes QwenVL’s performance on the XBD dataset. Figure 25 illustrates the capabilities enhanced by knowledge graphs.
Figure 25: XBD ResNet-EuroSAT Output Using QwenVL With Knowledge Graph
Image Caption [25] Caption: The image shows satellite view of town with mix of undamaged and damaged buildings there are undamaged buildings and damaged buildings visible in the scene the presence of both undamaged and damaged structures highlights the impact of the disaster on the affected area
The KG-enhanced model generates a succinct, precise evaluation that clearly delineates the mixed damage state—essential information for prioritizing rescue and recovery efforts.
In the absence of KG integration (Figure 26), the model’s output is verbose, redundant, and less dependable.
Figure 26: XBD ResNet-EuroSAT Output Using QwenVL Without Knowledge Graph
The baseline caption demonstrates phrase redundancy (“showing signs of destruction" is repeated thrice), incomplete phrases (“collapsed trees broken"), and formatting inaccuracies (“areasbuildings" concatenated without spacing). These problems diminish caption effectiveness and may obstruct automated processing systems.
Image Caption [26] Caption: the satellite image depicts an area that has been significantly impacted by tornado the visible damage is extensive with numerous buildings and structures showing signs of destruction showing signs of destruction here is detailed description of the visible damage and its implications damage overview damaged areas buildings these areas show visible signs of damage such as showing signs of destruction collapsed trees broken and debris scattered around the damage is concentrated in specific areas indicating that the tornado had localized impact undamaged areasbuildings these areas appear to be relatively intact with no visible signs of damage from the tornado
7.1 Summary of Findings
In all assessed configurations, KG integration uniformly enhances caption quality via three principal mechanisms:
  1. 1.
    Elimination of hallucinations: KG-enhanced models prevent the fabrication of details (casualty figures, fictitious individuals) that could mislead disaster response initiatives.
  2. 2.
    Improved coherence: Organized knowledge mitigates redundancy, fragmented sentences, and illogical word amalgamation that afflict baseline models.
  3. 3.
    Enhanced semantic accuracy: KG-grounded captions accurately delineate damage types, disaster categories, and structural conditions, offering actionable intelligence for emergency responders.
These enhancements are consistent across various model architectures (ViT, ResNet-EuroSAT), caption generators (LLaVA, QwenVL), and disaster datasets (RescueNet, XBD), illustrating the generalizability of KG-enhanced methodologies for disaster image captioning.
8 Conclusion
Image captioning involves assessing visual content and creating a textual description that highlights key parts of the image. Beyond object recognition, image captioning requires inferring information from unlabeled images. In this paper, we have presented a novel image captioning architecture. Our Vision Language Caption Enhancer (VLCE) framework is a more advanced framework for automated disaster image captioning. Our dual-architecture approach, combining both CNN-LSTM and transformer-based models with domain-specific backbones. It outperformed the baseline vision-language models and enriched knowledge across satellite and UAV images.
Future research should focus on generalization in low-resource domains, human-centric evaluation metrics, and lightweight architectures for real-world deployment. Techniques like knowledge graphs and VLMs, as demonstrated in our work, offer promising pathways for adaptable systems but require further testing in unstructured contexts (e.g., medical imaging). Furthermore, domain-specific applications require metrics that prioritize actionable insights over linguistic fluency, aligning with end-user needs in fields like disaster response. By combining VLMs, knowledge graphs, and reference-free evaluation, our contributions provide a blueprint for robust, domain-aware captioning systems, advancing the field toward scalable and interpretable solutions.